according to-S=ut+1/2at2 ,u=0
a=2S/t2

ACCORDING to definition-- a=v/t u=0
v=S/t
a=S/t2

WHY IS IT SO THAT BOTH a ARE DIFFERENT?

Grade:9

1 Answers

vikas askiitian expert
13 years ago

we use v = s/t when it is given that velocity is constant wrt time or no acclerated motion....

but when there is constant accleration then we use V = u + at  or s=ut+at2/2        (for constant accleration)